When a solution prepared by dissolving 4.00 g of an unknown monoprotic acid in 1.00 L of water is titrated with 0.600 M , 37.0 mL of the solution is needed to neutralize the acid.


a

Determine the molarity of the acid solution.


1
Expert's answer
2021-06-22T04:42:57-0400

1. Find the number of moles of 37ml (0.6M) solution used to neutralize the acid. To do this, use the following formula: n=V(l)*CM

n=V*CM=0.037L *0.6M=0.0222mol

2. 1 mole of alkali is used to neutralize 1 mole of monoprotic acid. that is, the monoprotic acid was also 0.0222mol

3. We use the following formula to find the molarity of an acid: CM=n/V

CM=0.0222mol/1L= 0.0222mol/L

Answer: 0.0222mol/L